OpenStudy (anonymous):

simplify 3x^2+17x-28/42-x-x^2

OpenStudy (anonymous):

3x^2+21x-4x-28/42-7x+6x-x^2 i hope now you can solve it.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you don't know how to solve quadratic equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is why i am here to get help

OpenStudy (anonymous):

ok i will solve it but you have to learn it.

OpenStudy (anonymous):

i just need to bring my grade up

OpenStudy (anonymous):

3x^2+21x-4x-28/42-7x+6x-x^2 = 3x(x+7)-4(x+7)/7(6-x)+x(6-x) =(x+7)(3x-4)/(6-x)(7+x) =3x-4/6-x

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x-4/6-x final answer.

OpenStudy (anonymous):

ok